### v1

@@ Abstract
The sequence form, owing to its compact and holistic strategy representation, has demonstrated significant efficiency in computing normal-form perfect equilibria for two-player extensive-form games with perfect recall. Nevertheless, the examination of $n$-player games remains underexplored. To tackle this challenge, we present a sequence-form characterization of normal-form perfect equilibria for $n$-player extensive-form games, achieved through a class of perturbed games formulated in sequence form. Based on this characterization, we develop a differentiable path-following method for computing normal-form perfect equilibria and prove its convergence. This method involves constructing an artificial logarithmic-barrier game in sequence form, where an additional variable is incorporated to regulate the influence of logarithmic-barrier terms to the payoff functions, as well as the transition of the strategy space. We prove the existence of a smooth equilibrium path defined by the artificial game, starting from an arbitrary positive realization plan and converging to a normal-form perfect equilibrium of the original game as the additional variable approaches zero. Furthermore, we extend Harsanyi’s linear and logarithmic tracing procedures to the sequence form and develop two alternative methods for computing normal-form perfect equilibria. Numerical experiments further substantiate the effectiveness and efficiency of our methods.
@@
Extensive-Form Game, Sequence Form, Normal-Form Perfect Equilibrium, Differentiable Path-Following Method
@@ 1 Introduction
Game theory provides a comprehensive mathematical framework for decision optimization in settings involving strategic interaction among rational agents. The primary concerns within this field revolve around the representation and resolution of games. The extensive-form game [1] provides a significant representation, especially applicable in scenarios involving sequential interactions. As a central solution concept for extensive-form games, Nash equilibrium signifies a state in which no player can improve their payoff by changing their strategy alone. Nevertheless, as explored by Selten [2], Myerson [3], Kreps and Wilson [4], and van Damme [5], an extensive-form game can have multiple Nash equilibria, some of which may deviate from our intuitive expectations regarding the game’s outcome and off-equilibrium strategies. As a refinement of Nash equilibrium, perfect equilibrium introduced by Selten [2] can eliminate a variety of counter-intuitive equilibria by introducing slight perturbations to the strategies. According to Selten, the concept of perfect equilibrium in extensive-form games can be classified into two types: extensive-form perfect equilibrium and normal-form perfect equilibrium, neither of which is contained within the other. As noted by Van Damme [6], extensive-form perfect equilibria may involve dominated strategies. In addition, Kohlberg and Mertens [7] contend that the reduced normal form contains all essential information required for decision-making. Stalnaker [8] further emphasizes the sufficiency of normal-form representations in the context of epistemic models. Therefore, it is crucial to study the equilibrium refinements of extensive-form games in normal form. This paper focuses on the computation of normal-form perfect equilibria for finite $n$-player extensive-form games with perfect recall.
@@
The typical methods for computing normal-form perfect equilibria in extensive-form games rely on transforming these games into their normal-form representations, followed by the application of equilibrium computation methods specific to normal-form games. Notably, path-following methods have gained prominence as powerful and effective tools for computing equilibria in normal-form games. These methods are fundamentally grounded in the computation of Nash equilibria. Lemke and Howson [9] proposed a complementarity pivoting algorithm to obtain Nash equilibria for bimatrix games. Subsequent extensions independently by Rosenmüller [10] and Wilson [11] adapted this algorithm to $n$-player games. To render this extended method computationally feasible, Garcia et al. [12] introduced a simplicial path-following method and implemented it to approximate Nash equilibria. Over the subsequent decades, significant advancements were made in the development of simplicial path-following methods for computing Nash equilibria in normal-form games. Noteworthy contributions include the work of van der Laan and Talman [13], Doup and Talman [14], and Herings and van den Elzen [15], who proposed increasingly flexible and efficient methods. Although these methods are capable of ultimately reaching Nash equilibria, their convergence rates are notably hindered by the failure to exploit the differentiability inherent in games. In response, several differentiable path-following methods have been presented in the literature, with key contributions from Herings and Peeters [16], Harsanyi and Selten  [17], Govindan and Wilson [18], and Chen and Dang [19]. These approaches have significantly enhance the convergence rates for computing Nash equilibria in normal-form games.
@@
Research into the computation of Nash equilibria in normal-form games has led to the development of various methods for determining perfect equilibria. van den Elzen and Talman [20] developed the first method to compute a perfect equilibrium, employing a complementary pivoting algorithm that operates exclusively in bimatrix games. Chen and Dang [21] generalized Kohlberg and Mertens’s Nash equilibrium reformulation [7, 19] to perturbed games and proposed a simplicial path-following method for identifying a perfect equilibrium in $n$-person games. Subsequently, Chen and Dang [22] extended the logistic quantal response equilibrium, demonstrating the existence of a smooth path to a perfect equilibrium under a specific assumption regarding payoff functions. Chen and Dang [23] introduced an exterior-point differentiable homotopy method capable of selecting an approximate perfect equilibrium. Motivated by the limitations of prior methods and guided by the principles of Harsanyi and Selten’s equilibrium selection philosophy [17], more alternative schemes for computing perfect equilibria have been developed. By exploiting the selection properties inherent in the Nash’s mappings, Harsanyi’s tracing procedures, and logistic quantal response equilibrium, Cao and Dang [24, 25, 26] developed their respective variants and distinct differentiable path-following methods to select an exact perfect equilibrium in normal-form games. Differentiable path-following methods have also demonstrated excellent performance in computing other refinements of Nash equilibria [27, 28].
@@
While various differentiable path-following methods for computing perfect equilibria in normal-form games exist, the exponential growth in the size of normal-form representations severely hampers computational efficiency, making such methods infeasible for even medium-scale games. Despite the considerable contraction in size accomplished through reduced normal-form representations, the issue of exponential growth remains persistent [29]. To mitigate the complexity arising from this transformation, Wilson [30] and Koller and Megiddo [31] suggested using mixed strategies with small supports in two-player extensive-form games. Koller and Megiddo [32] introduced a polynomial-time algorithm that applies the concept of realization weight on nodes to solve two-person zero-sum extensive-form games with perfect recall, where the number of constraints increases exponentially. von Stengel [33] later proposed a more compact strategy representation, known as the sequence form, thereby reformulating the Nash equilibrium computation in a more computationally efficient manner. In this representation, pure strategies are replaced by sequences and random strategies comply with a recursively defined linear system. Building upon the sequence form, notable progress has been achieved for computing equilibria in two-player extensive-form games with perfect recall. Koller et al. [34] developed an algorithm for identifying Nash equilibria by applying Lemke’s algorithm to the linear complementarity problem derived from the sequence-form representation. This algorithm’s efficiency was experimentally validated through the Gala system developed by Koller and Pfeffer [35]. von Stengel et al. [36] extended the method proposed by van den Elzen and Talman [20] for computing perfect equilibria in two-player normal-form games to the sequence form, enabling the computation of normal-form perfect equilibria in extensive-form games. Miltersen and Sørensen [37] modified the algorithm proposed by Koller et al. [34] to accommodate perturbed games, facilitating the computation of quasi-perfect equilibria. Studies on game situations with $n$ players remain limited. Govindan and Wilson [38] extended structure theorems to perturbed extensive-form games using enabling strategies, which, in essence, mirror strategies in sequence form. This extension brought about a piecewise differentiable path-following method for computing Nash equilibria in $n$-player perturbed extensive-form games. Furthermore, no globally differentiable path-following methods have been proposed to support the computation of normal-form perfect equilibria. A significant challenge in the advancement of computational methods arises from the absence of theoretical foundations for the characterization of normal-form perfect equilibria in sequence form. Although Gatti et al. [39] introduced a sequence-form characterization of quasi-perfect equilibria, their approach relies on a sequential structure. This sequential dependency conflicts with the simultaneity principle of the normal form, rendering it inadequate for fully characterizing all normal-form perfect equilibria.
@@
The objective of this study is to develop a sequence-form characterization of normal-form perfect equilibria in $n$-player extensive-form games with perfect recall and, based on this, to propose an effective and efficient globally differentiable path-following method for their computation. To achieve this, we begin by establishing the equivalence relationship between strategies in normal-form and sequence-form representations, as well as their connection to best-response strategies. Following this, we introduce a class of perturbed games in sequence form and apply an optimization-based approach to derive the corresponding equilibrium systems for these perturbed games. We then demonstrate the necessity and sufficiency of the solution limits of these equilibrium systems for characterizing normal-form perfect equilibria. To develop a differentiable path-following method based on this characterization, we construct an artificial game in sequence form by introducing an additional variable and incorporating logarithmic barrier terms into the payoff functions. As the additional variable decreases from two to zero, the artificial game undergoes two distinct stages of transformation. The first stage serves to locate a unique starting point, where strategies transitioned from constructed strategies to realization plans. In the second stage, the logarithmic-barrier terms restrict the strategy to the interior of the realization plan space, thereby ensuring a well-defined transformation from realization plans to mixed strategies. We establish the existence of a smooth equilibrium path dictated by the artificial game, which, as the additional variable tends towards zero, converges to a normal-form perfect equilibrium of the original game. Finally, we conduct numerical experiments to validate the effectiveness and efficiency of our proposed methods.
@@
The remaining sections of this paper are organized as follows. Section 2 provides a review of normal-form perfect equilibrium in extensive-form games and the sequence form. In Section 3, we present a sequence-form characterization of normal-form perfect equilibria. In Section 4, we propose a sequence-form globally differentiable logarithmic path-following method to compute normal-form perfect equilibria. In Section 5, we extend Harsanyi’s linear and logarithmic tracing procedures to the sequence form and develop two alternative computational methods. The numerical performance and comparative analysis of these methods are reported in Section 6, and the paper concludes with Section 7.
2 Notation and Preliminaries
@@
$H$ Set of histories, $\emptyset\in H$ and $\langle a_{1},\ldots,a_{L}\rangle\in H$ if $\langle a_{1},\ldots,a_{K}\rangle\in H$ and $L<K$
@@
$I^{j}_{i}\in\mathcal{I}_{i},j\in M_{i}$ $j$th information set of player $i\in N_{c}$, $A(I^{j}_{i})\triangleq A(h)=A(h^{\prime})$ whenever $h,h^{\prime}\in I^{j}_{i}$
@@
The notation and conventions for extensive-form games are adopted from Osborne and Rubinstein [40] and outlined in Table 1. An extensive-form game is represented by
@@
In this paper, Our focus is on finite extensive-form games with perfect recall. “finite” means that $H$ is a finite set. Perfect recall holds if, for each player $i$, any histories $h$ and $h^{\prime}$ in the same information set satisfy $R_{i}(h)=R_{i}(h^{\prime})$, ensuring consistent memory of past actions and knowledge.
@@
The chance player’s mixed strategy $\sigma^{c}=(\sigma^{c}(s^{c}):s^{c}\in S^{c})$ is fixed and determined by $\sigma^{c}(s^{c})=\prod_{h\in H,P(h)=c}\sum_{a\in A(h)}s^{c}(a)f_{c}(a|h)$. Additional notations and their descriptions are provided in Table 2. Then the normal-form representation of $\Gamma$ is expressed as $\Gamma_{n}=\langle N,S,\sigma^{c},\{u^{i}\}_{i\in N}\rangle$.
In the reduced normal-form representation, pure strategies are defined in a more compact manner while preserving all valid strategic information. Specifically, for a pure strategy $s^{i}$ of player $i\in N_{c}$, $s^{i}(I^{j}_{i})=a$, $j\in M_{i},a\in A(I^{j}_{i})$ means that, for $\langle a_{1},\ldots,a_{L}\rangle\in I^{j}_{i}$, $s^{i}(I^{j_{q}}_{i})=a_{q}$ holds for all $0\leq q\leq L$ with $j_{q}\in M_{i},a_{q}\in A(I^{j_{q}}_{i})$. All other definitions remain unchanged and are still applicable. To highlight the superiority of our methods, all derivations in this paper are based on the reduced normal form. For simplicity, we shall refer to it as the normal form throughout, omitting the qualifier “reduced”.
@@
A mixed strategy profile $\sigma^{*}$ is referred to a Nash equilibrium if, for every player $i\in N$, the inequality $u^{i}(\sigma^{*})\geq u^{i}(\sigma^{i},\sigma^{*-i})$ is satisfied for all $\sigma^{i}\in\Xi^{i}$. This condition ensures that no player can improve their payoff by unilaterally deviating from their strategy in the equilibrium profile. However, the weakness of this condition can lead to a large equilibrium set, leading to the emergence of numerous counterintuitive equilibria and great uncertainty in determining which equilibrium to choose. In response to this limitation, Selten [2] introduced the concept of perfect equilibrium, eliminating a large number of unreasonable equilibria. The definition of normal-form perfect equilibrium in an extensive-form game is as follows.
@@
$w^{i}$ Sequence of actions taken by player $i$
$w^{i}_{I^{j}_{i}}$ Sequence of player $i$ leading to $I^{j}_{i}$, $w^{i}_{h}=w^{i}_{I^{j}_{i}}$ for any $h\in I^{j}_{i}$
$w^{i}_{I^{j}_{i}}a$ The extended sequence $w^{i}_{I^{j}_{i}}\cup\{a\}$
${W}=\underset{i\in N_{c}}{\times}{W}^{i}$ The collection of sequence profiles, $\emptyset\in{W}^{i}$
$g^{i}(w)$ Expected payoff of player $i$ on the sequence profile $w$
@@
$M_{i}(w^{i})$ The index set of the information sets for player $i$ with $w^{i}$ being the sequence
$D_{i}$ The set of $(j,a)$ for player $i$ with $M_{i}(w^{i}_{I^{j}_{i}}a)=\emptyset$
@@
The computation of a normal-form perfect equilibrium of an extensive-form game typically requires a transformation into its normal form. As Wilson [30] points out, even simple extensive-form games often produce exceedingly large normal forms due to the exponential increase in the number of pure strategies relative to the number of information sets. To circumvent this exponential growth, the sequence form, formally developed by von Stengel [33], has emerged as a particularly efficient alternative.
@@
The sequence form replaces pure strategies with sequences, providing a compact representation. For $i\in N_{c}$, a sequence $w^{i}$ is defined as the action set of player $i$ for some history. Specifically, for $h=\langle a_{1},\ldots,a_{L}\rangle\in H$, the corresponding sequence is given by
$w^{i}_{h}=\{a_{q}:\text{$a_{q}\in A(I^{j}_{i})$ for some $j\in M_{i}$ and $1\leq q\leq L$}\},$
@@
which is either the empty sequence $\emptyset$ or an extension $w^{i}_{I^{j}_{i}}a$ of a preceding sequence $w^{i}_{I^{j}_{i}}$ with $i\in N$, $j\in M_{i}$, and $a\in A(I^{j}_{i})$. The function $g^{i}$ determines the payoff for player $i$ in any sequence profile $w\in{W}$, defined as
@@
$g^{i}(w)=\left\{\begin{array}[]{ll}u^{i}_{z}(h)&\text{if $w$ is defined by $h\in Z$,}\\ 0&\text{otherwise.}\end{array}\right.$
We say that $w=(w^{i}:i\in N_{c})\in{W}$ is defined by $h=\langle a_{1},\ldots,a_{L}\rangle$ if $\mathop{\cup}_{i\in N_{c}}w^{i}=\{a_{1},\ldots,a_{L}\}$.
@@
The inherent challenge in developing algorithms for the sequence form lies in the fact that randomization over sequences is no longer represented by a straightforward probability distribution. Instead, it requires the formulation of an recursive system of linear equations. For player $i\in N_{c}$, a random strategy in the sequence form is a function $\gamma^{i}$ defined on ${W}^{i}$, with the convention that $\gamma^{i}(\emptyset)=1$. We call $\gamma^{i}$ a realization plan for player $i$ if it satisfies the linear system,
$\begin{array}[]{l}\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(w^{i}_{I^{j}_{i}}a)-\gamma^{i}(w^{i}_{I^{j}_{i}})=0,\;j\in M_{i},\\ 0\leq\gamma^{i}(w^{i}_{I^{j}_{i}}a),\;j\in M_{i},a\in A(I^{j}_{i}).\end{array}$ (3)

### v3

@@ Abstract
The sequence form, owing to its compact and holistic strategy representation, has demonstrated significant efficiency in computing normal-form perfect equilibria for two-player extensive-form games with perfect recall. Nevertheless, the examination of $n$-player games remains underexplored. To tackle this challenge, we present a sequence-form characterization of normal-form perfect equilibria for $n$-player extensive-form games, achieved through a class of perturbed games formulated in sequence form. Based on this characterization, we develop a differentiable path-following method for computing normal-form perfect equilibria and prove its convergence. This method formulates an artificial logarithmic-barrier game in sequence form, introducing an additional variable to regulate the impact of logarithmic-barrier terms on the payoff functions, as well as the transition of the strategy space. We prove the existence of a smooth equilibrium path defined by the artificial game, starting from an arbitrary positive realization plan and converging to a normal-form perfect equilibrium of the original game as the additional variable approaches zero. Furthermore, we extend Harsanyi’s linear and logarithmic tracing procedures to the sequence form and develop two alternative methods for computing normal-form perfect equilibria. Numerical experiments further substantiate the effectiveness and computational efficiency of our methods.
@@
[…]
@@ 1 Introduction
Game theory provides a comprehensive mathematical framework for decision optimization in settings involving strategic interaction among rational agents. The primary concerns within this field revolve around the representation and resolution of games. The extensive-form game [KuhnExtensiveGames1950] provides a significant representation, especially applicable in scenarios involving sequential interactions. As a central solution concept for extensive-form games, Nash equilibrium signifies a state in which no player can improve their payoff by changing their strategy alone. Nevertheless, as explored by Selten [SeltenReexaminationperfectnessconcept1975], Myerson [MyersonRefinementsNashequilibrium1978], Kreps and Wilson [KrepsSequentialEquilibria1982], and van Damme [vanDammeStabilityperfectionNash1987], an extensive-form game can have multiple Nash equilibria, some of which may deviate from our intuitive expectations regarding the game’s outcome and off-equilibrium strategies. As a refinement of Nash equilibrium, perfect equilibrium introduced by Selten [SeltenReexaminationperfectnessconcept1975] can eliminate a variety of counter-intuitive equilibria by introducing slight perturbations to the strategies. According to Selten, the concept of perfect equilibrium in extensive-form games can be classified into two types: extensive-form perfect equilibrium and normal-form perfect equilibrium, neither of which is contained within the other. As noted by van Damme [vandammeRelationPerfectEquilibria1984], extensive-form perfect equilibria may involve dominated strategies. In addition, Kohlberg and Mertens [KohlbergStrategicStabilityEquilibria1986] contended that the reduced normal form contains all essential information required for decision-making. Stalnaker [stalnakerExtensiveStrategicForms1999] further emphasized the sufficiency of normal-form representations in the context of epistemic models. Therefore, it is crucial to study the equilibrium refinements of extensive-form games in normal form. This paper investigates the computation of normal-form perfect equilibria in finite $n$-player extensive-form games with perfect recall.
@@
The typical methods for computing normal-form perfect equilibria in extensive-form games rely on transforming these games into their normal-form representations, followed by the application of equilibrium computation methods specific to normal-form games. Notably, path-following methods have gained prominence as powerful and effective tools for computing equilibria in normal-form games. These methods are fundamentally grounded in the computation of Nash equilibria. Lemke and Howson [LemkeEquilibriumPointsBimatrix1964] proposed a complementarity pivoting algorithm to obtain Nash equilibria for bimatrix games, and later, various other algorithms in [vandenelzenProcedureFindingNash1991, aliExploringTwoNew2024, hussainNumericalExplorationTwo2025] for bimatrix games have been developed. Subsequent extensions by Rosenmüller [RosenmullerGeneralizationLemkeHowson1971] and Wilson [WilsonComputingEquilibriaNPerson1971] independently adapted the Lemke-Howson algorithm to $n$-player games. To render this extended method computationally feasible, Garcia et al. [GarciaSimplicialApproximationEquilibrium1973] introduced a simplicial path-following method and implemented it to approximate Nash equilibria. Over the subsequent decades, significant advancements were made in the development of simplicial path-following methods for computing Nash equilibria in normal-form games. Noteworthy contributions include the work of van der Laan and Talman [vanderLaanComputationFixedPoints1982], Doup and Talman [Doupnewsimplicialvariable1987], and Herings and van den Elzen [HeringsComputationNashEquilibrium2002], who proposed increasingly flexible and efficient methods. Although these methods are capable of ultimately reaching Nash equilibria, their convergence rates are notably hindered by the failure to exploit the differentiability inherent in games [allgowerPiecewiseLinearMethods2000a]. In response, several differentiable path-following methods have been presented in the literature, with key contributions from Herings and Peeters [Heringsdifferentiablehomotopycompute2001], Harsanyi and Selten  [HarsanyiGeneralTheoryEquilibrium1988], Govindan and Wilson [GovindanglobalNewtonmethod2003], and Chen and Dang [Chenreformulationbasedsmoothpathfollowing2016]. These approaches have markedly enhanced the convergence rates when computing Nash equilibria in normal-form games.
@@
Research into the computation of Nash equilibria in normal-form games has led to the development of various methods for determining perfect equilibria. van den Elzen and Talman [vandenelzenProcedureFindingNash1991] developed the first method to compute a perfect equilibrium, employing a complementary pivoting algorithm that operates exclusively in bimatrix games. Chen and Dang [chenReformulationBasedSimplicialHomotopy2019] generalized Kohlberg and Mertens’s Nash equilibrium reformulation [KohlbergStrategicStabilityEquilibria1986, Chenreformulationbasedsmoothpathfollowing2016] to perturbed games and proposed a simplicial path-following method for identifying a perfect equilibrium in $n$-person games. Subsequently, Chen and Dang [Chenextensionquantalresponse2020] extended the logistic quantal response equilibrium, demonstrating the existence of a smooth path to a perfect equilibrium under a specific assumption regarding payoff functions. Chen and Dang [Chendifferentiablehomotopymethod2021] introduced an exterior-point differentiable homotopy method capable of selecting an approximate perfect equilibrium. Motivated by the limitations of prior methods and guided by the principles of Harsanyi and Selten’s equilibrium selection philosophy [HarsanyiGeneralTheoryEquilibrium1988], more alternative schemes for computing perfect equilibria have been developed. By exploiting the selection properties inherent in the Nash’s mappings, Harsanyi’s tracing procedures, and logistic quantal response equilibrium, Cao and Dang [caoComplementarityEnhancedNashs2022, CaovariantHarsanyitracing2022, caoVariantLogisticQuantal2024] developed their respective variants and distinct differentiable path-following methods to select an exact perfect equilibrium in normal-form games. Differentiable path-following methods have also shown remarkable performance in computing other refinements of Nash equilibria [Zhansmoothpathfollowingalgorithm2018, caoDifferentiablePathFollowingMethod2023].
@@
While several differentiable path-following methods exist for computing perfect equilibria in normal-form games, the exponential growth inherent in normal-form representations significantly undermines computational tractability, rendering these methods impractical even for games of moderate scale. Although reduced normal-form representations achieve notable size reductions, the fundamental problem of exponential expansion persists [dalkey1953equivalence]. To mitigate the complexity arising from this transformation, Wilson [wilsonComputingEquilibriaTwoPerson1972] and Koller and Megiddo [KollerFindingmixedstrategies1996] suggested employing mixed strategies with small supports in two-player extensive-form games. Furthermore, Koller and Megiddo [Kollercomplexitytwopersonzerosum1992] developed a polynomial-time algorithm that uses realization weights at nodes to solve two-person zero-sum extensive-form games with perfect recall, despite the exponential increase in constraints. To overcome these scalability issues, von Stengel [vonStengelEfficientComputationBehavior1996] introduced the sequence form, a more compact representation that reformulates Nash equilibrium computation with improved efficiency. In this representation, pure strategies are replaced by sequences and random strategies comply with a recursively defined linear system. Building upon this, substantial progress has been made for two-player extensive-form games with perfect recall. Koller et al. [KollerEfficientComputationEquilibria1996] developed an algorithm that applies Lemke’s method [LemkeBimatrixEquilibriumPoints1965] to the linear complementarity problem derived from the sequence form. Its practical efficiency was demonstrated through the Gala system by Koller and Pfeffer [KollerRepresentationssolutionsgametheoretic1997]. Subsequently, von Stengel et al. [VonStengelComputingNormalForm2002] extended van den Elzen and Talman’s [vandenelzenProcedureFindingNash1991] method for computing perfect equilibria in normal-form games to the sequence form, enabling the computation of normal-form perfect equilibria in extensive-form games. Miltersen and Sørensen [MiltersenComputingquasiperfectequilibrium2010] further adapted Koller et al.’s [KollerEfficientComputationEquilibria1996] approach to handle perturbed games, facilitating the computation of quasi-perfect equilibria. Research on games involving more than two players remains limited. Govindan and Wilson [govindanStructureTheoremsGame2002] extended structure theorems for perturbed extensive-form games by introducing strategies in sequence form, leading to a piecewise differentiable path-following method for computing Nash equilibria in $n$-player perturbed extensive-form games. Hou et al. [houSequenceformDifferentiablePathfollowing2025] proposed a globally differentiable path-following method within the sequence form framework to compute Nash equilibria. However, no sequence-form globally differentiable path-following methods currently exist to compute normal-form equilibrium refinements, including normal-form perfect equilibria. A major barrier to further progress is the lack of a theoretical foundation for characterizing normal-form perfect equilibria in sequence form. Although Gatti et al. [gattiCharacterizationQuasiperfectEquilibria2020] presented a sequence-form characterization for quasi-perfect equilibria, their reliance on sequential structures conflicts with the simultaneity principle inherent in the normal form, making it insufficient for fully capturing normal-form perfect equilibria.
@@
The objective of this study is to develop a sequence-form characterization of normal-form perfect equilibrium in $n$-player extensive-form games with perfect recall and, based on this, to propose an effective and efficient globally differentiable path-following method for its computation. To achieve this, we begin by establishing the equivalence relationship between strategies in normal-form and sequence-form representations, as well as their connection to best-response strategies. Following this, we introduce a class of perturbed games in sequence form and apply an optimization-based approach to derive the corresponding equilibrium systems for these perturbed games. We then demonstrate the necessity and sufficiency of the solution limits of these equilibrium systems for characterizing normal-form perfect equilibria. To develop a differentiable path-following method based on this characterization, we construct an artificial game in sequence form by introducing an additional variable and embedding logarithmic barrier terms within the payoff functions. As the additional variable decreases from two to zero, the artificial game experiences two distinct transformation phases. The initial phase aims to identify a unique starting point by shifting strategies from the constructed form to realization plans. In the subsequent phase, the logarithmic barrier terms confine the strategies within the interior of the realization plan space, thus guaranteeing a well-defined mapping from realization plans to mixed strategies. We establish the existence of a smooth equilibrium path governed by the artificial game, which converges to a normal-form perfect equilibrium of the original game as the additional variable tends towards zero. Lastly, numerical experiments are performed to demonstrate the effectiveness and computational efficiency of our proposed methods.
@@
The rest of this paper is structured as follows. Section 2 reviews the concept of normal-form perfect equilibrium within extensive-form games and introduces the sequence form. In Section 3, we provide a sequence-form characterization of normal-form perfect equilibria. Section 4 proposes a globally differentiable logarithmic path-following method based on the sequence form for computing normal-form perfect equilibria. In Section 5, we extend Harsanyi’s linear and logarithmic tracing procedures to the sequence form, resulting in two alternative computational approaches. Numerical results and comparative evaluations of these methods are presented in Section 6, with concluding remarks given in Section 7.
2 Preliminaries
@@
$H$ Set of histories,
$\emptyset\in H$ and $\langle a_{1},\ldots,a_{L}\rangle\in H$ if $\langle a_{1},\ldots,a_{K}\rangle\in H$ and $L<K$
@@
$I^{j}_{i}\in\mathcal{I}_{i},j\in M_{i}$ $j$th information set of player $i\in N_{c}$,
$A(I^{j}_{i})\triangleq A(h)=A(h^{\prime})$ whenever $h,h^{\prime}\in I^{j}_{i}$
@@
We adopt the notation for extensive-form games from Osborne and Rubinstein [OsborneCourseGameTheory1994], summarized in Table 1. An extensive-form game can be represented by
@@
This paper addresses finite extensive-form games with perfect recall. “finite" means that $H$ is a finite set. Perfect recall holds if, for each player $i$, any histories $h$ and $h^{\prime}$ in the same information set satisfy $R_{i}(h)=R_{i}(h^{\prime})$, ensuring consistent memory of past actions and knowledge.
@@
The chance player’s mixed strategy $\sigma^{c}=(\sigma^{c}(s^{c}):s^{c}\in S^{c})$ is fixed and determined by $\sigma^{c}(s^{c})=\prod_{h\in H,P(h)=c}\sum_{a\in A(h)}s^{c}(a)f_{c}(a|h)$. Additional notations and their descriptions are provided in Table 2. Then the normal-form representation of $\Gamma$ is expressed as $\Gamma_{n}=\langle N,S,\Xi,\{u^{i}\}_{i\in N}\rangle$.
In the reduced normal-form representation, pure strategies are defined in a more compact manner while preserving all valid strategic information. Specifically, for a pure strategy $s^{i}$ of player $i\in N_{c}$, $s^{i}(I^{j}_{i})=a$, $j\in M_{i},a\in A(I^{j}_{i})$ means that, for $\langle a_{1},\ldots,a_{L}\rangle\in I^{j}_{i}$, $s^{i}(I^{j_{q}}_{i})=a_{q}$ holds for all $0\leq q\leq L$ with $j_{q}\in M_{i},a_{q}\in A(I^{j_{q}}_{i})$. All other definitions remain unchanged and are still applicable. To highlight the superiority of our methods, all derivations in this paper are based on the reduced normal form. For simplicity, we shall refer to it as the normal form throughout, omitting the qualifier “reduced".
@@
A mixed strategy profile $\sigma^{*}$ is referred to a Nash equilibrium if, for every player $i\in N$, the inequality $u^{i}(\sigma^{*})\geq u^{i}(\sigma^{i},\sigma^{*-i})$ is satisfied for all $\sigma^{i}\in\Xi^{i}$. This condition ensures that no player can improve their payoff by unilaterally deviating from their strategy in the equilibrium profile. However, the weakness of this condition can lead to a large equilibrium set, leading to the emergence of numerous counterintuitive equilibria and great uncertainty in determining which equilibrium to choose. In response to this limitation, Selten [SeltenReexaminationperfectnessconcept1975] introduced the concept of perfect equilibrium, eliminating a large number of unreasonable equilibria. The definition of normal-form perfect equilibrium in an extensive-form game is as follows.
@@
$\varpi^{i}$ Sequence of actions taken by player $i$
$\varpi^{i}_{I^{j}_{i}}$ Sequence of player $i$ leading to $I^{j}_{i}$, $\varpi^{i}_{h}=\varpi^{i}_{I^{j}_{i}}$ for any $h\in I^{j}_{i}$
$\varpi^{i}_{I^{j}_{i}}a$ Extended sequence $\varpi^{i}_{I^{j}_{i}}\cup\{a\}$
${W}=\underset{i\in N_{c}}{\times}{W}^{i}$ Collection of sequence profiles, $\emptyset\in{W}^{i}$
$g^{i}(\varpi)$ Expected payoff of player $i$ on the sequence profile $\varpi$
@@
$M_{i}(\varpi^{i})$ Set of $j\in M_{i}$ satisfying $\varpi^{i}_{I^{j}_{i}}=\varpi^{i}$
$D_{i}$ Set of $(j,a)$ for player $i$ with $M_{i}(\varpi^{i}_{I^{j}_{i}}a)=\emptyset$
@@
The computation of a normal-form perfect equilibrium of an extensive-form game typically requires a transformation into its normal form. As Wilson [wilsonComputingEquilibriaTwoPerson1972] points out, even simple extensive-form games often produce exceedingly large normal forms due to the exponential increase in the number of pure strategies relative to the number of information sets. To circumvent this exponential growth, the sequence form, formally developed by von Stengel [vonStengelEfficientComputationBehavior1996], has emerged as a particularly efficient alternative.
@@
The sequence form replaces pure strategies with sequences, providing a compact representation. For $i\in N_{c}$, a sequence $\varpi^{i}$ is defined as the action set of player $i$ for some history. Specifically, for $h=\langle a_{1},\ldots,a_{L}\rangle\in H$, the corresponding sequence is given by
$\varpi^{i}_{h}=\{a_{q}:\text{$a_{q}\in A(I^{j}_{i})$ for some $j\in M_{i}$ and $1\leq q\leq L$}\},$
@@
which is either an empty sequence $\emptyset$ or an extension $\varpi^{i}_{h^{\prime}}a$ of a preceding sequence $\varpi^{i}_{h^{\prime}}$ with $i\in N$ and $h^{\prime}\in H$. The function $g^{i}$ specifies the payoff of player player $i$ for any sequence profile $\varpi\in{W}$, defined by
@@
$g^{i}(\varpi)=\left\{\begin{array}[]{ll}u^{i}_{z}(h)&\text{if $\varpi$ is defined by $h\in Z$,}\\ 0&\text{otherwise.}\end{array}\right$.
We say that $\varpi=(\varpi^{i}:i\in N_{c})\in{W}$ is defined by $h=\langle a_{1},\ldots,a_{L}\rangle$ if $\mathop{\cup}_{i\in N_{c}}\varpi^{i}=\{a_{1},\ldots,a_{L}\}$.
@@
The inherent challenge in developing algorithms for the sequence form lies in the fact that randomization over sequences is no longer represented by a probability distribution over the strategies at an information set. Instead, it requires the formulation of an recursive system of linear equations. For player $i\in N_{c}$, a random strategy in the sequence form is a function $\gamma^{i}$ defined on ${W}^{i}$, with the convention that $\gamma^{i}(\emptyset)=1$. We call $\gamma^{i}$ a realization plan for player $i$ if it satisfies the linear system,
$\begin{array}[]{l}\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(\varpi^{i}_{I^{j}_{i}}a)-\gamma^{i}(\varpi^{i}_{I^{j}_{i}})=0,\;j\in M_{i},\\ 0\leq\gamma^{i}(\varpi^{i}_{I^{j}_{i}}a),\;j\in M_{i},a\in A(I^{j}_{i}).\end{array}$ (3)